# Localization of One-Dimensional Random Band Matrices

### v1

@@ Abstract.
We consider a general class of $n\times n$ random band matrices with bandwidth $W.$ When $W^{2}\ll n$, we prove that with high probability the eigenvectors of such matrices are localized and decay exponentially at the sharp scale $W^{2}$. Combined with the delocalization results of Erdős and Riabov [21], and Yau and Yin [42], this establishes the conjectured localization-delocalization transition for a large class of random band matrices.
@@ 1.1. Model and Results
(2) $\left\|\frac{d^{2}}{dx^{2}}\left(\log\phi\right)\left(x\right)\right\|_{L^{\infty}}\leq M.$
@@
Combined with the delocalization results of Erdős and Riabov [21], Theorem 1 establishes the conjectured localization-delocalization transition for a large class of one-dimensional RBMs. Specifically, let $H$ be as in Theorem 1 and additionally suppose the variances in each row sum to $1$, and the entries of $\sqrt{W}H$ have uniformly bounded moments of all orders. If $W^{2}/n\gg 1$, Corollary 3.5 of [21] implies that all eigenvectors of $H$ with eigenvalues in $2+\epsilon,2-\epsilon]$ are delocalized with high probability. Conversely, if $W^{2}/n\ll 1$, Theorem 1 implies that all such eigenvectors of $H$ are exponentially localized with high probability. We note that [21] generalizes the work of Yau and Yin [42] on the classical Wegner orbital model.
@@ 1.3. Prior Mathematical Work
The first rigorous mathematical progress on the localization side of Conjecture 1 was due to Schenker [32]. He proved localization in a large class of $n\times n$ Gaussian RBMs when $W^{8}/n\ll 1$, and for a wider class of non-Gaussian models he proved there exists a $C>0$ such that localization holds when $W^{C}/n\ll 1$.
Subsequently, Schenker’s argument was refined by several groups of authors. Using a new sharp Wegner estimate for Gaussian matrices, Peled, Schenker, Shamis, and Sodin [31] proved localization in a class of Gaussian RBMs when $W^{7}/n\ll 1$. Later, this was improved to $W^{4}/n\ll 1$ in the real Wegner orbital model, which is obtained by taking $A_{i}$ and $B_{i}$ to be GOE and Ginibre matrices, respectively. This was done contemporaneously by Chen and Smart [16], and Cipolloni, Peled, Schenker, and Shapiro [18] (whose proof extends to Gaussian-type models). Most recently, a preprint of Goldstein [28] claims to prove localization in the real Wegner orbital model when $N\geq W^{2}$, i.e. $W^{3}/n\ll 1$, with eigenfunctions localized at the essentially sharp scale $(\log W)^{3}W^{2}$. It is possible that the ideas in [28] may be sharpened to reach the regime $W^{2}/n\ll 1$ in that model; however, to our knowledge, the community has been unable to verify the proof, and shortly after posting [28], Goldstein passed away.
@@
The delocalization side of Conjecture 1 was recently resolved. In the classical Wegner orbital model, i.e. $A_{i}$ and $B_{i}$ are GUE and complex Ginibre matrices, the work of Yau and Yin [42] proved delocalization of eigenvectors, GUE-type eigenvalue statistics, and QUE when $W^{2}/n\gg 1$. The work of Erdős and Riabov [21] extended these results to a very general class of band matrices. These two works are part of a recent series of spectacular breakthroughs on delocalization in RBMs. For instance see [19, 41, 20, 23].
@@ 1.4. Proof Sketch
where $v_{j}$ is as in (6). Both choices of sign in (10) vary $\alpha_{j}$ by $2W^{-1/2}$ and maintain the direction of $B_{j}v_{j}$. Hence, if we show at least one choice of sign in (10) decreases the density of $(B_{j}\ |\mathcal{F})$ by at most an $O(1)$ factor, we expect $\alpha_{j}$ to fluctuate at the scale $W^{-1/2}$. This argument is abstractly formulated in Lemma 2. The change in the density of $\left(B_{j}|\mathcal{F}\right)$ under (10) is essentially computed by Taylor expansion, and the quantities $\left\|B_{j}v_{j}\right\|,\left\|A_{j+1}v_{j}\right\|,$ and $\left\|B_{j}^{\ast}D_{j}^{-1}B_{j}v_{j}\right\|$ arise naturally from this computation. See Lemma 3 for the details.
@@ 3. Fluctuations Lower Bound
The idea is to replace $B_{j}$ with $|B_{j}\pm\delta\ket{B_{j}v_{j}}\bra{v_{j}}$ and compute the distortion of the density of $B_{j}$. Either choice of sign changes $\alpha_{j}$ by $\delta$, and so if one of the choices usually decreases the density of $B_{j}$ by at most an $O(1)$ factor, then $\alpha_{j}$ should fluctuate at the scale $\delta$. This argument is made rigorous in the following Mermin–Wagner type lemma. On first reading it is instructive to imagine $X$ is a Gaussian on $\mathbb{R}$, $T_{\pm}(x)=\left(1\pm\delta\right)x$, and $g(x)=\log\left|x\right|$.

### v2

@@ Abstract.
We consider a general class of $n\times n$ random band matrices with bandwidth $W.$ When $W^{2}\ll n$, we prove that with high probability the eigenvectors of such matrices are localized and decay exponentially at the sharp scale $W^{2}$. Combined with the delocalization results of Yau and Yin [44], and Erdős and Riabov [22], this establishes the conjectured localization-delocalization transition for a large class of random band matrices.
@@ 1.1. Model and Results
(2) $\left\|\frac{d^{2}}{dx^{2}}\left(\log\phi\right)\left(x\right)\right\|_{L^{\infty}(\mathbb{R})}\leq M.$
@@
Combined with the delocalization results in [44] and [22], Theorem 1 establishes the conjectured localization-delocalization transition for a large class of one-dimensional RBMs. Specifically, let $H$ be as in Theorem 1 and additionally suppose the variances in each row sum to $1$, and the entries of $\sqrt{W}H$ have uniformly bounded moments of all orders. If $W^{2}/n\gg 1$, Corollary 3.5 of [22] implies that all eigenvectors of $H$ with eigenvalues in $2+\epsilon,2-\epsilon]$ are delocalized with high probability. Conversely, if $W^{2}/n\ll 1$, Theorem 1 implies that all such eigenvectors of $H$ are exponentially localized with high probability.
@@ 1.3. Prior Mathematical Work
The first rigorous mathematical progress on the localization side of Conjecture 1 was due to Schenker [34]. He proved localization in a large class of $n\times n$ Gaussian RBMs when $W^{8}/n\ll 1$, and for a wider class of non-Gaussian models he proved there exists a $C>0$ such that localization holds when $W^{C}/n\ll 1$. The argument, explained below, combines a fluctuation estimate for $G(1,N)$ with an a-priori estimate, and was first proposed by Michael Aizenman.
Subsequently, Schenker’s argument was refined by several groups of authors. Using a new sharp Wegner estimate for Gaussian matrices, Peled, Schenker, Shamis, and Sodin [32] proved localization in a class of Gaussian RBMs when $W^{7}/n\ll 1$. Later, this was improved to $W^{4}/n\ll 1$ in the real Wegner orbital model, which is obtained by taking $A_{i}$ and $B_{i}$ to be GOE and Ginibre matrices, respectively. This was done contemporaneously by Chen and Smart [16], and Cipolloni, Peled, Schenker, and Shapiro [18] (whose proof extends to mixtures of Gaussian models). Most recently, a preprint of Goldstein [29] claims to prove localization in the real Wegner orbital model when $N\geq W^{2}$, i.e. $W^{3}/n\ll 1$, with eigenfunctions localized at the essentially sharp scale $(\log W)^{3}W^{2}$. It is possible that the ideas in [29] may be sharpened to reach the regime $W^{2}/n\ll 1$ in that model; however, to our knowledge, the community has been unable to verify the proof, and shortly after posting [29], Goldstein passed away.
@@
The delocalization side of Conjecture 1 was recently resolved. In the classical Wegner orbital model, i.e. $A_{i}$ and $B_{i}$ are GUE and complex Ginibre matrices, Yau and Yin [44] proved delocalization of eigenvectors, GUE-type eigenvalue statistics, and QUE when $W^{2}/n\gg 1$. The work of Erdős and Riabov [22] extended these results to a very general class of band matrices. These two works are part of a recent series of spectacular breakthroughs on delocalization in RBMs. For instance see [20, 43, 21, 24].
@@ 1.4. Proof Sketch
where $v_{j}$ is as in (6). Both choices of sign in (10) vary $\alpha_{j}$ by $2W^{-1/2}$ and maintain the direction of $B_{j}v_{j}$. Hence, if we show at least one choice of sign in (10) decreases the density of $(B_{j}\ |\mathcal{F})$ by at most an $O(1)$ factor, we expect $\alpha_{j}$ to fluctuate at the scale $W^{-1/2}$. This argument is abstractly formulated in Lemma 2. The change in the density of $\left(B_{j}|\mathcal{F}\right)$ under (10) is estimated by Taylor expansion, and the quantities $\left\|B_{j}v_{j}\right\|,\left\|A_{j+1}v_{j}\right\|,$ and $\left\|B_{j}^{\ast}D_{j}^{-1}B_{j}v_{j}\right\|$ arise naturally from this computation. See Lemma 3 for the details. The proof of Lemma 2 is related to Pfister’s [33] and Dobrushin and Shlosman’s [19] proof of the Mermin-Wagner Theorem in Statistical Mechanics, and we note this idea was first introduced to RBMs in [18].
@@ 3. Fluctuations Lower Bound
The idea is to replace $B_{j}$ with $|B_{j}\pm\delta\ket{B_{j}v_{j}}\bra{v_{j}}$ and compute the distortion of the density of $B_{j}$. Either choice of sign changes $\alpha_{j}$ by $\delta$, and so if one of the choices usually decreases the density of $B_{j}$ by at most an $O(1)$ factor, then $\alpha_{j}$ should fluctuate at the scale $\delta$. This argument is made rigorous in the following Mermin–Wagner type lemma. The work [18] was the first to lower bound fluctuations in this setting using a Mermin-Wagner argument. On first reading it is instructive to imagine $X$ is a Gaussian on $\mathbb{R}$, $T_{\pm}(x)=\left(1\pm\delta\right)x$, and $g(x)=\log\left|x\right|$.